# Psychiatry-Bench: A Multi-Task Benchmark for LLMs in Psychiatry

### v1

@@ 7.1 Limitations
The foremost limitation is a direct consequence of the critical research gap we identify: the absence of large-scale, publicly available, and expert-validated clinical datasets for mental health. Forced a reliance on meticulous manual curation from authoritative textbooks. While this process ensures unparalleled clinical validity, the core contribution of our work is inherently labor-intensive. As a result, the scale of PsychiatryBench, though substantial, is smaller than benchmarks built on synthetic data or automated web-scraping. It was a deliberate trade-off to prioritize clinical safety over sheer data volume.
@@
Furthermore, our reliance on published textbooks introduces a trade-off between authoritative knowledge and real-world clinical practice. The vignettes in these resources are, by design, well-structured and pedagogically focused to illustrate specific diagnostic principles. They may not fully capture the ambiguity, noise, and unstructured nature of raw clinical data, such as real-time patient dialogues or fragmented electronic health record (EHR) notes. Consequently, while PsychiatryBench robustly tests an LLM’s reasoning against expert-curated knowledge, performance may not perfectly generalize to the complexities of live, unfiltered clinical interactions.
@@
A related point is that the distribution of psychiatric disorders within PsychiatryBench does not mirror their real-world epidemiological prevalence. It is a direct consequence of our source materials; psychiatric textbooks often over-represent rare, complex, or "classic" cases to serve specific pedagogical goals. Therefore, models evaluated on PsychiatryBench are tested on their ability to handle a teaching-oriented distribution of cases, which differs from the case mix they would encounter in a general clinical setting.
@@
Finally, our evaluation methodology for open-ended tasks relies on LLM-as-a-judge. While this approach offers the scalability and consistency required to evaluate thousands of responses across numerous models, it is not a perfect substitute for human expert evaluation. LLM judges can have inherent biases and may not capture the full clinical nuance apparent to a practicing psychiatrist. We mitigated this by grounding the evaluation in expert-validated reference answers, meaning the judge’s task was one of comparing an output to a "gold standard" rather than making a de novo clinical judgment. However, a full-scale validation by a panel of human psychiatrists was beyond the scope of this work and remains a crucial future direction.
@@ 7.2 Future Work
Our analysis identifies seven key priorities for advancing LLM applications in mental health, summarized in the following points:
  • Expanded Dataset Coverage: Future research should aim to include a wider range of psychiatric disorders, cultural contexts, and multimodal data sources (e.g., EHRs, imaging, and audio transcripts) to create a more representative benchmark. Collaborations with international psychiatric organizations could facilitate access to diverse clinical data.
  • Automated and Scalable Annotation: Developing robust automated tools for case extraction and annotation, validated by expert psychiatrists, would enable larger-scale datasets while maintaining clinical accuracy. Techniques like natural language processing (NLP) for structured data extraction could streamline this process.
  • Multimodal Evaluation Frameworks: Integrating multimodal inputs, such as lab results or patient audio, into evaluation datasets would better simulate real-world clinical decision-making. Could involve adapting LLMs to process structured data alongside text or developing hybrid models for multimodal reasoning.
  • Broader LLM Comparisons: Evaluating a diverse set of LLMs, including open-source models and those fine-tuned for medical domains, would provide deeper insights into their relative strengths and weaknesses in psychiatric applications.
  • Enhanced Evaluation Metrics: Future work should focus on designing metrics that prioritize clinical safety, cultural sensitivity, and robustness against hallucinations. Incorporating human-in-the-loop evaluation by practicing psychiatrists could ensure alignment with clinical standards.
  • Comprehensive Task Diversity: Expanding the task set to include empathetic dialogue, crisis intervention, and longitudinal care planning would enable a more complete assessment of LLMs’ potential in mental healthcare. Could involve creating conversational benchmarks like MentalChat 16K but grounded in expert-validated data.
  • Prompt Optimization Frameworks: Systematic exploration of the prompt space, using automated prompt engineering tools, could identify optimal prompt structures for clinical tasks. Standardizing prompt formats across studies would improve reproducibility and comparability of results.
@@ 8 Conclusion
In this work, we have introduced PsychiatryBench, the first comprehensive, clinically grounded benchmark for evaluating LLMs in psychiatric applications. By curating over 5,300 question-answer pairs from eleven authoritative textbooks and casebooks, ranging from diagnostic vignettes and treatment planning to longitudinal follow-up and complex sequential reasoning, we ensure that model performance is measured against expert-validated, real-world clinical knowledge rather than unverified or synthetic data. Our eleven distinct task types probe the full spectrum of psychiatric reasoning, from factual recall and classification to multi-turn decision-making and management planning.
Through rigorous evaluation of both general-purpose and medical-specialized LLMs, including proprietary systems such as Gemini 2.5 Pro and MedGemma, as well as open-source models like LLaMA and Palmyra-Med, we identify clear performance leaders and persistent gaps. Frontier models achieve strong alignment on single-turn tasks such as diagnosis and management planning (up to 8̃7 % similarity). However, performance degrades on multi-step follow-up, classification of comorbid disorders, and prompt robustness across varying formats. Our novel similarity scoring, along with consistency metrics for full versus separated EMIs and multi-label classification, provides a nuanced picture of models’ strengths and limitations in high-stakes clinical assessments.
These findings underscore both the promise and the perils of deploying LLMs in mental healthcare. On one hand, leading models demonstrate impressive capacity for clinical reasoning, empathetic response, and knowledge synthesis; on the other, the risk of hallucinations, prompt sensitivity, and difficulties with multi-label, longitudinal scenarios highlight the necessity of human oversight and continued refinement.
Looking ahead, PsychiatryBench offers a modular framework for future research. Extending the dataset with additional case complexities, such as cross-cultural presentations, pediatric and forensic psychiatry, and integrating multimodal inputs (e.g., speech, imaging) will further bridge the gap between AI evaluation and real-world clinical practice. Moreover, fine-tuning and retrieval-augmented approaches hold promise for improving model robustness and safety. We encourage the community to leverage PsychiatryBench as both a diagnostic tool to benchmark new models and as a scaffolding for developing clinically reliable, interpretable AI systems that can augment psychiatric care while rigorously safeguarding patient well-being.

### v2

@@ 6.1 Limitations
6.1.1 Data Source and Textbook Limitations
@@
The foremost limitation is that our reliance on published textbooks introduces a trade-off between authoritative knowledge and real-world clinical practice. The vignettes in these resources are, by design, well-structured and pedagogically focused to illustrate specific diagnostic principles. They may not fully capture the ambiguity, noise, and unstructured nature of raw clinical data, such as real-time patient dialogues or fragmented electronic health record (EHR) notes. Consequently, while PsychiatryBench robustly tests an LLM’s reasoning against expert-curated knowledge, performance may not perfectly generalize to the complexities of live, unfiltered clinical interactions.
@@
A related point is that the distribution of psychiatric disorders within PsychiatryBench does not mirror their real-world epidemiological prevalence. This is a direct consequence of our source materials; psychiatric textbooks often over-represent rare, complex, or "classic" cases to serve specific pedagogical goals. Therefore, models evaluated on PsychiatryBench are tested on their ability to handle a teaching-oriented distribution of cases, which differs from the case mix they would encounter in a general clinical setting. Additionally, PsychiatryBench’s case distribution emphasizes moderate-acuity outpatient presentations, with limited representation of acute crisis scenarios such as severe psychosis requiring emergency intervention, active suicidal intent with plan and means, or manic episodes necessitating involuntary hospitalization.
6.1.2 Western Psychiatric Bias
A significant limitation of PsychiatryBench is its reliance on predominantly Western psychiatric sources, particularly American textbooks such as DSM-5-TR Clinical Cases and Stahl’s Essential Psychopharmacology. While these resources represent gold-standard, expert-validated knowledge within Western psychiatric practice, they inherently reflect Western conceptualizations of mental illness, symptom presentation, diagnostic categories, and treatment approaches. Psychiatric presentations, the validity of diagnostic constructs, and the appropriateness of treatment recommendations can vary significantly across cultures. Help-seeking behaviors, symptom expression, and therapeutic expectations are all influenced by cultural context, yet our current benchmark evaluates LLMs primarily on their alignment with Western psychiatric frameworks. Consequently, strong performance on PsychiatryBench does not necessarily indicate an LLM’s readiness for deployment in culturally diverse or non-Western clinical settings.
6.1.3 Evaluation Methodology Limitations
@@
Our evaluation methodology for open-ended tasks relies on an LLM-as-a-judge framework. While this approach offers the scalability and consistency required to evaluate thousands of responses across numerous models, it is not a perfect substitute for human expert evaluation. LLM judges can exhibit inherent biases and may not capture the full clinical nuance apparent to a practicing psychiatrist. A significant methodological limitation concerns the potential circularity of our approach: the model selected as the evaluator (Llama 3 70B) was also included as one of the models under evaluation in our test set. We mitigated this by grounding the evaluation in expert-validated reference answers, meaning the judge’s task was one of comparing outputs to a “gold standard” rather than making de novo clinical judgments. However, a full-scale validation by a panel of human psychiatrists was beyond the scope of this work and remains a crucial future direction. LLM-as-judge evaluation also introduces potential bias due to overlapping reasoning frameworks. In particular, when Llama 3 70B serves as both evaluator and evaluated model, a direct circularity arises that may artificially inflate agreement metrics, representing an acknowledged methodological limitation. Strong performance on the expert-validated psychiatric knowledge and reasoning tasks used by PsychiatryBench does not necessarily indicate clinical preparedness. Additional validation, such as prospective clinical trials, evaluations of conversational coherence in patient contacts, assessments of cultural competence across varied groups, and strict safety monitoring in real-world clinical settings, would be necessary for real-world clinical deployment.
A further limitation is the unavoidable risk of training data contamination. The authoritative, publicly available textbooks and casebooks used to construct PsychiatryBench are likely included in the pre-training corpora for many of the large-scale models evaluated. This may inflate performance scores, particularly on tasks measuring factual recall or knowledge recognition. We believe the benchmark’s value is nonetheless retained, as its primary focus is on novel task formats that test the application of knowledge and multi-step clinical reasoning rather than simple rote memorization.
6.1.4 Safety Limitations
The current evaluation framework does not systematically assess several safety-critical behaviors essential for the responsible use of LLMs in psychiatric contexts. Models may inadvertently reinforce delusional beliefs by favoring conversational coherence over clinical accuracy, thereby exacerbating pathological thinking. For instance, LLMs may also mishandle suicidal ideation by failing to recognize acute risk or escalate appropriately, a shortcoming with potentially life-threatening consequences. Additional concerns include the tendency toward anthropomorphism using language such as “I’m here for you,” which can blur therapeutic boundaries and foster unhealthy attachments, as well as sycophantic responses that prioritize user satisfaction over evidence-based reasoning, potentially minimizing symptoms or validating unsafe preferences. These behaviors reflect forms of value misalignment that extend far beyond factual correctness and are not captured by our current LLM-as-a-judge scoring approach, which focuses primarily on semantic agreement with expert references. As a result, clinicians and ethicists have raised concerns that reliance on opaque, uncalibrated systems may erode clinical judgment, amplify existing biases, and undermine fairness in mental healthcare, underscoring the urgent need for evaluation frameworks that directly interrogate safety, alignment, and boundary-maintaining behavior rather than indirect proxies of correctness.
6.1.5 Additional Limitations
Additional limitations include: the variable representation of medical mimics (e.g., thyroid disease, delirium, CNS pathology) that was not systematically designed to test for medical rule-outs; the intentional limitation to adult psychiatry, excluding child and adolescent psychiatry which represents a distinct subspecialty; and the primary focus on diagnostic reasoning and treatment planning, with limited assessment of broader integrative skills such as comprehensive biopsychosocial formulation, functional assessments, or multidisciplinary collaboration planning.
@@ 6.2 Future Work
Future work should build directly on the limitations identified above and focus on strengthening the clinical realism, cultural breadth, methodological rigor, and safety of LLM evaluation in psychiatry. These directions aim to advance the development of more comprehensive, reliable, and clinically defensible AI systems that reflect the complexities of real-world psychiatric practice.
6.2.1 Expanding Clinical Data Representativeness
Addressing the limitations associated with textbook-derived vignettes requires the incorporation of more diverse and ecologically valid clinical data sources. Future versions of PsychiatryBench should integrate de-identified electronic health records, naturalistic patient–clinician dialogues, and real-time clinical notes that capture the ambiguity, noise, and fragmentation characteristic of authentic psychiatric encounters. This expansion would allow for evaluation settings that more closely approximate the unpredictability and variability of real-world practice. Furthermore, the dataset should broaden its diagnostic coverage to better align with real-world epidemiology by including a larger proportion of highly prevalent disorders, as well as more medically complex presentations. Enriching the dataset with emergency and crisis-level scenarios such as active suicidal intent, severe psychosis, or acute mania requiring involuntary intervention would further enhance ecological validity. Scalable data curation pipelines that combine automated extraction tools with expert psychiatric oversight would facilitate the responsible and accurate expansion of clinically representative cases.
6.2.2 Improving Cross-Cultural Validity
To address the Western psychiatric bias inherent in the current benchmark, future work should incorporate case materials and diagnostic frameworks from a broader range of cultural and geographic contexts. Integrating guidelines and casebooks from Asia, Africa, Latin America, and the Middle East would promote greater cultural representativeness and reduce reliance on Western-centric models of mental illness. Expanding the benchmark to include culturally variable idioms of distress, distinct help-seeking patterns, and diverse symptom presentations would help assess whether models can avoid applying ethnocentric assumptions or misinterpreting culturally mediated clinical information. Collaborations with international psychiatric organizations and culturally diverse clinical institutions will be essential for acquiring representative source material. Additionally, the development of evaluation metrics that explicitly assess cultural competence would allow for systematic examination of whether models accurately and sensitively interpret psychiatric presentations across varying cultural contexts.
6.2.3 Improving Evaluation Methodology and Reducing Model Circularity
Future methodological work should focus on strengthening the reliability and validity of LLM-as-a-judge evaluation frameworks. Incorporating panels of licensed psychiatrists for periodic benchmarking would provide a clinical anchor point for evaluating whether automated scoring methods align with expert judgment. Establishing adjudication systems that combine human oversight with LLM-based evaluators may further reduce the biases associated with relying on a single model for judgment. To mitigate circularity, future iterations of the benchmark should ensure strict separation between the models serving as evaluators and those being evaluated, potentially through the development of judge models specifically designed for assessment tasks. Enhancing the evaluation methodology to include step-by-step reasoning audits, diagnostic chain-of-thought analyses, and temporal generalization tests based on post–knowledge cutoff data would offer a more robust assessment of clinical reasoning capabilities. Further research is also needed to quantify and mitigate the effects of training data contamination, including temporal ablation analyses and the use of synthetic or newly developed case vignettes that can test true generalization rather than memorization.
6.2.4 Safety-Critical Evaluation and Value Alignment
Advancing the safety evaluation of LLMs is essential for their use in psychiatric contexts. Future work should develop comprehensive evaluation tasks that explicitly test models’ handling of clinically sensitive scenarios. These tasks must assess whether models can avoid reinforcing delusional content, appropriately recognize and respond to suicidal ideation, maintain clear therapeutic boundaries without anthropomorphic self-representation, and prioritize clinical accuracy over user agreement in situations where user preferences conflict with evidence-based practice. Developing evaluation metrics that systematically examine these behaviors will enable a more rigorous assessment of safety vulnerabilities that extend beyond factual correctness. Long-term directions should include the integration of clinically informed red-teaming efforts, formal ethical oversight, and the creation of value alignment benchmarks that test whether models can recognize the limits of their competence, decline unsafe tasks, and default to risk-minimizing behaviors during ambiguous or high-risk situations. Strengthening safety evaluation in this manner will be critical for ensuring that LLMs support rather than undermine clinician judgment and patient well-being.
6.2.5 Broadening Clinical Scope and Integrative Task Coverage
Future expansions of PsychiatryBench should extend its clinical breadth to encompass domains currently underrepresented or absent. Incorporating cases involving medical mimics such as endocrinological disorders, neurological disease, and delirium would enable evaluation of integrated medical–psychiatric reasoning, which is essential in many real-world settings. Further work should introduce a comprehensive suite of cases from child and adolescent psychiatry, a subspecialty with distinct diagnostic frameworks, developmental considerations, and contextual factors. Beyond diagnosis and treatment planning, the benchmark should also evaluate broader integrative capabilities, including biopsychosocial formulation, functional assessment, and the capacity to incorporate psychosocial stressors into clinical reasoning. Evaluating longitudinal care abilities, such as tracking symptom evolution, anticipating relapse risk, and adjusting treatment plans over time, would align more closely with the realities of ongoing psychiatric care. Incorporating these dimensions will allow future versions of PsychiatryBench to capture the full complexity of psychiatric practice and provide a more comprehensive foundation for evaluating LLMs intended for clinical use.
@@ 7 Conclusion
PsychiatryBench represents a significant step toward establishing clinically grounded, domain-specific standards for evaluating LLMs in mental healthcare. By integrating thousands of expert-curated QA pairs derived from authoritative psychiatric texts, alongside complex multi-step clinical vignettes, the benchmark captures a broad spectrum of diagnostic, therapeutic, and longitudinal reasoning skills. Our analysis reveals that advanced models can demonstrate strong psychiatric knowledge and generate coherent clinical interpretations, yet these strengths are unevenly distributed across task types. Models continue to struggle with ambiguity, multi-label diagnostic reasoning, context-dependent interpretation, and scenarios requiring high-fidelity clinical judgment. These limitations highlight the risks of deploying general-purpose models in sensitive psychiatric settings without rigorous, domain-tailored evaluation. PsychiatryBench thus fills a critical gap by providing a structured, transparent, and reproducible framework for assessing the depth and reliability of psychiatric reasoning in contemporary LLMs.
At the same time, this benchmark serves as a foundation for future innovations in AI-driven mental healthcare. Improving the cultural breadth, ecological validity, and modality diversity of PsychiatricBench will be essential to creating more comprehensive assessments that reflect the realities of global clinical practice. Equally important will be the development of safety-oriented evaluations that specifically test crisis response, boundary maintenance, hallucination robustness, and value alignment dimensions that are central to real-world psychiatric care but absent from most current benchmarks. Continued interdisciplinary collaboration between psychiatrists, clinical researchers, computational scientists, and ethicists will be necessary to ensure that future models not only perform well on structured tasks but also adhere to the ethical and safety standards expected in mental health contexts. By establishing a strong evaluative foundation and outlining clear directions for progress, PsychiatryBench aims to catalyze the creation of AI systems that meaningfully support clinicians, enhance patient care, and advance the safe and equitable integration of AI in psychiatric settings.